If 72% of 25 students like maths, find out the number of students who do not like mathematics?

Solution:

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

72% of 25 students like maths.

Thus, 72% of 25 = (72/100) × 25 = 18 students

Now, from 25 students, 18 students like maths

Hence, the number of students who do not like maths = 25 – 18 = 7 students

Thanks